# State-space kinetic Ising model reveals task-dependent entropy flow in sparsely active nonequilibrium neuronal dynamics

### v2

@@ Methods
${\sigma}_{t}^{\rm forward}$ $\simeq-\sum_{\mathbf{x}_{t-1}}Q(\mathbf{x}_{t-1})\sum_{\mathbf{x}_{t}}p(\mathbf{x}_{t}|\mathbf{x}_{t-1})\log p(\mathbf{x}_{t}|\mathbf{x}_{t-1})$
$=-\sum_{\mathbf{x}_{t-1}}Q(\mathbf{x}_{t-1})\sum_{i}\sum_{x_{i,t}}p(x_{i,t}|\mathbf{x}_{t-1})\log p(x_{i,t}|\mathbf{x}_{t-1})$
$=\sum_{\mathbf{x}_{t-1}}Q(\mathbf{x}_{t-1})\sum_{i}\chi\left(h_{i,t}(\mathbf{x}_{t-1})\right),$ (46)
@@
$\phantom{====}-(1-r(h_{i,t}(\mathbf{x}_{t-1})))\log(1-r(h_{i,t}(\mathbf{x}_{t-1})))$
$=-\left(r(h_{i,t}(\mathbf{x}_{t-1})\log\frac{r(h_{i,t}(\mathbf{x}_{t-1}))}{1-r(h_{i,t}(\mathbf{x}_{t-1}))}-\log(1-r(h_{i,t}(\mathbf{x}_{t-1}))\right)$ (47)
Furthermore, by defining $\Gamma_{i,t}$,
$\Gamma_{i,t}$ $=\sum_{\mathbf{x}_{t-1}}Q(\mathbf{x}_{t-1})\chi\left(h_{i,t}(\mathbf{x}_{t-1})\right),$ (48)
${\sigma}^{\rm forward}_{t}$ can be summarized as
${\sigma}^{\rm forward}_{t}$ $=\sum_{i}\Gamma_{i,t}.$ (49)
Here, we approximate $\Gamma_{i,t}$ by a Gaussian distribution based on the central limit theorem for a collection of independent binary signals. Specifically, by using $\mathcal{D}_{z}=\frac{\mathrm{d}z}{\sqrt{2\pi}}\exp\left(-\frac{1}{2}z^{2}\right)$, $\Gamma_{i,t}$ is approximated as
$\Gamma_{i,t}$ $=\sum_{\mathbf{x}_{t-1}}Q\left(\mathbf{x}_{t-1}\right)\chi\left(h_{i,t}(\mathbf{x}_{t-1})\right)\approx\int\mathcal{D}_{z}\,\chi\left(g_{i,t}+z\sqrt{\Delta_{i,t}}\right)$ (50)
where $g_{i,t}$ and $\Delta_{i,t}$ are the mean of $h_{i,t}(\mathbf{x}_{t-1})$ given by
$g_{i,t}$ $\equiv\theta_{i,t}+\sum_{j}\theta_{ij,t}m_{j,t-1},$ (51)
$\Delta_{i,t}$ $\equiv\operatorname{Var}\left[h_{i,t}\right]=\sum_{j}\theta_{ij,t}^{2}m_{j,t-1}(1-m_{j,t-1}).$ (52)
Here, $m_{i,t}$ ($t=1,\ldots,T$) is the mean-field approximation of $x_{i,t}$ obtained by the Gaussian approximation method assuming independent activity of neurons at t1t-1:
$m_{i,t}$ $\equiv\sum_{\mathbf{x}_{t},\mathbf{x}_{t-1}}x_{i,t}p(\mathbf{x}_{t}|\mathbf{x}_{t-1})p(\mathbf{x}_{t-1})$
$\simeq\sum_{\mathbf{x}_{t},\mathbf{x}_{t-1}}x_{i,t}p(\mathbf{x}_{t}|\mathbf{x}_{t-1})Q(\mathbf{x}_{t-1})$
$=\sum_{\mathbf{x}_{t-1}}Q(\mathbf{x}_{t-1})\sum_{\mathbf{x}_{t}}x_{i,t}p(\mathbf{x}_{t}|\mathbf{x}_{t-1})$
$=\sum_{\mathbf{x}_{t-1}}Q(\mathbf{x}_{t-1})r(h_{i,t}(\mathbf{x}_{t-1}))$ (53)
@@
$m_{i,t}$ $\simeq\sum_{\mathbf{x}_{t-1}}r(h_{i,t}(\mathbf{x}_{t-1}))Q(\mathbf{x}_{t-1})\approx\int\mathcal{D}_{z}\,r\left(g_{i,t}+z\sqrt{\Delta_{i,t}}\right),$ (54)
for $t=1,\ldots,T$, using Eqs. 51 and 52, which are functions of $m_{i,t-1}$. Here $m_{i,1}$ was computed using nominal values of $m_{i,0}$ ($i=1,\ldots,N$). In the simulation and empirical analyses, we used spiking probability averaged over all time steps and trials for each neuron as $m_{i,0}$.
@@
$=-\sum_{\mathbf{x}_{t-2}}p(\mathbf{x}_{t-2})\sum_{\mathbf{x}_{t},\mathbf{x}_{t-1}}p(\mathbf{x}_{t}|\mathbf{x}_{t-1})p(\mathbf{x}_{t-1}|\mathbf{x}_{t-2})\log{p(\mathbf{x}_{t-1}|\mathbf{x}_{t})}$
$=-\sum_{\mathbf{x}_{t},\mathbf{x}_{t-1}}\sum_{\mathbf{x}_{t-2}}p(\mathbf{x}_{t}|\mathbf{x}_{t-1})p(\mathbf{x}_{t-1}|\mathbf{x}_{t-2})p(\mathbf{x}_{t-2})\sum_{i}\left[x_{i,t-1}h_{i,t}(\mathbf{x}_{t})-\psi(h_{i,t}(\mathbf{x}_{t}))\right]$
$=-\sum_{\mathbf{x}_{t-2}}\sum_{\mathbf{x}_{t-1}}p(\mathbf{x}_{t-1}|\mathbf{x}_{t-2})p(\mathbf{x}_{t-2})\sum_{\mathbf{x}_{t}}p(\mathbf{x}_{t}|\mathbf{x}_{t-1})\sum_{i}\left[x_{i,t-1}h_{i,t}(\mathbf{x}_{t})-\psi(h_{i,t}(\mathbf{x}_{t}))\right],$ (55)
where we redefine $\psi$ as a function of $h_{i,t}(\mathbf{x}_{t})$: $\psi(h_{i,t}(\mathbf{x}_{t}))=\log(1+e^{h_{i,t}(\mathbf{x}_{t})})$. We also note that $\psi(h_{i,t}(\mathbf{x}_{t}))=-\log(1-r(h_{i,t}(\mathbf{x}_{t}))$.
We note that
$\sum_{\mathbf{x}_{t-1}}p(\mathbf{x}_{t-1}|\mathbf{x}_{t-2})$ $=\sum_{\mathbf{x}_{t-1}}\prod_{i}p(x_{i,t-1}|\mathbf{x}_{t-2})$
$=\sum_{x_{1,t-1}}\sum_{x_{2,t-1}}\cdots\sum_{x_{N,t-1}}\prod_{i}p(x_{i,t-1}|\mathbf{x}_{t-2})$
$=\prod_{i}\sum_{x_{i,t-1}}p(x_{i,t-1}|\mathbf{x}_{t-2}).$ (56)
We approximate the following probabilities by independent distributions.
@@
Using these equations, ${\sigma}^{\rm backward}_{t}$ can be approximated as follows:
${\sigma}^{\rm backward}_{t}$ $\simeq-\sum_{\mathbf{x}_{t-2}}\sum_{\mathbf{x}_{t-1}}p(\mathbf{x}_{t-1}|\mathbf{x}_{t-2})Q(\mathbf{x}_{t-2})\sum_{\mathbf{x}_{t}}Q(\mathbf{x}_{t})\sum_{i}\left[x_{i,t-1}h_{i,t}(\mathbf{x}_{t})-\psi(h_{i,t}(\mathbf{x}_{t}))\right]$
$=-\sum_{\mathbf{x}_{t-2}}Q(\mathbf{x}_{t-2})\sum_{\mathbf{x}_{t}}Q(\mathbf{x}_{t})\sum_{\mathbf{x}_{t-1}}p(\mathbf{x}_{t-1}|\mathbf{x}_{t-2})\sum_{i}\left[x_{i,t-1}h_{i,t}(\mathbf{x}_{t})-\psi(h_{i,t}(\mathbf{x}_{t}))\right]$
$=-\sum_{\mathbf{x}_{t}}Q(\mathbf{x}_{t})\sum_{\mathbf{x}_{t-2}}Q(\mathbf{x}_{t-2})\sum_{i}\left[r(h_{i,t-1}(\mathbf{x}_{t-2}))h_{i,t}(\mathbf{x}_{t})-\psi(h_{i,t}(\mathbf{x}_{t}))\right]$
@@
where we used Eqs. 43, 56 to obtain the third equality and Eq. 53 to obtain the last result. By defining
$\phi_{i}(h_{i,t}(\mathbf{x}_{t}))=-(m_{i,t-1}h_{i,t}(\mathbf{x}_{t})-\psi(h_{i,t}(\mathbf{x}_{t}))),$ (60)
@@
${\sigma}^{\rm backward}_{t}$ $=\sum_{i}\sum_{\mathbf{x}_{t}}Q(\mathbf{x}_{t})\phi_{i}(h_{i,t}(\mathbf{x}_{t}))$
$\approx\sum_{i}\int\mathcal{D}_{z}\,\phi_{i}\left(\tilde{g}_{i,t}+z\sqrt{\tilde{\Delta}_{i,t}}\right),$ (61)
@@
$\tilde{g}_{i,t}$ $\equiv\theta_{i,t}+\sum_{j}\theta_{ij,t}m_{j,t},$ (62)
$\tilde{\Delta}_{i,t}$ $=\sum_{j}\theta_{ij,t}^{2}m_{j,t}(1-m_{j,t}).$ (63)
An alternative approach to obtain the backward conditional entropy is given in S2 text.
@@
$\approx\sum_{i}\int\mathcal{D}_{z}\,\left[-\chi\left(g_{i,t}+z\sqrt{\Delta_{i,t}}\right)+\phi_{i}\left(\tilde{g}_{i,t}+z\sqrt{\tilde{\Delta}_{i,t}}\right)\right],$ (64)
@@
See also Supporting Information 3 text for the analytical expression of the entropy flow under specific conditions.
@@
We generated time-varying field and coupling parameters which are used to generate the binary data and subsequently estimated from the data. To generate smooth temporal variations, each coupling parameter $\theta_{ij,t}$ was sampled from a Gaussian process of size $T$ with mean $\mu$ and covariance matrix defined by the squared exponential kernel
$k(t,t^{\prime})=k_{0}\exp\left({-\frac{(t-t^{\prime})^{2}}{2\tau^{2}}}\right).$ (65)
For the analysis of estimation error and computational time using different system sizes (Fig. 3), we used the scaling mean $\mu=\frac{5}{N}$ and variance $k_{0}=10/N$, following the convention of the Sherrington-Kirkpatrick model. The characteristic length-scale was specified by $\tau=\frac{30}{\sqrt{N}}$. Similarly, the external field parameters $\theta_{i,t}$ were independently sampled from the Gaussian process, using $\mu=-3$, $\tau=50$, and $k_{0}=1$.
To obtain the smooth trajectories with respect to the different system sizes, a single set of random values was generated for the maximum system size ($N_{\max}$), and subsets of these values were used to examine the system size $N$. Specifically, for the coupling parameters, a global three-dimensional array was created with dimensions corresponding to the maximum number of neurons, time steps, and coupling connections. Similarly, for the field parameters, a two-dimensional array was generated, with dimensions corresponding to the maximum number of time steps and neurons. For a given neuron count $N$, the relevant subset of values was extracted from these precomputed arrays, ensuring that each $N$ used a subset of the values assigned to larger $N$. This hierarchical structure ensured that the seed for $N=80$ encompassed all values used for smaller $N$, maintaining consistency across different system sizes. We evaluated the model’s performance using this data set and repeated the procedure 10 times.
@@
To perform the analysis on fitting the kinetic Ising model to a mismatched model, we generated binary spike sequences generated by the nonlinearity that goes beyond linear synaptic summation and a logistic activation function, therefore induces the higher-order interactions (HOIs) in the population activity, by employing the recently proposed alternating-shrinking HOI model[58].
@@
$p(\mathbf{x})=\frac{h\left(\sum_{i=1}^{N}x_{i}\right)}{Z}\exp\left[-f\sum_{j=1}^{N}\left(-1\right)^{j+1}C_{j}\left(\frac{\sum_{i=1}^{N}x_{i}}{N}\right)^{j}\right],$ (66)
where $f$ is a sparsity parameter and $Z$ is the partition function. $h\left(\sum_{i=1}^{N}x_{i}\right)$ is an entropy canceling base measure function defined using binomial coefficient:
$h\left(n\right)=1\left/\binom{N}{n}\right$. (67)
The parameters $C_{1},C_{2},\ldots,C_{N}$ are the shrinking parameters, where $C_{j}=\left(\tau\right)^{j}$ with  $0<\tau<1$ results in the shifted-geometric exponential distribution.
The population spike-count distribution is the probability distribution of $n$ active neurons in the binary patterns, i.e., $n=\sum_{i=1}^{N}x_{i}$, which is given as
$P\left(n\right)=$ $\binom{N}{n}\mathcal{P}\left(x_{1}=1,\ldots,x_{n}=1,x_{n+1}=0,\ldots,x_{N}=0\;|\;\boldsymbol{\theta}_{N}\right)$
@@
We performed Gibbs sampling from this distribution, which corresponds to recurrent neural networks with a threshold–supralinear activation nonlinearity. For neuron $i$, let $n=\sum_{j\neq i}x_{j}$ be the spike count of the other units, and define
$Q(n)$ $=\sum_{j=1}^{N}(-1)^{j+1}C_{j}\left(\frac{n}{N}\right)^{j},$ (69)
$\Delta Q(n)$ $=Q(n{+}1)-Q(n).$ (70)
The unnormalized conditional weights are
$p_{0}$ $\propto h(n)\,\exp\!\bigl(-f\,Q(n)\bigr)\quad(x_{i}=0),$ (71)
$p_{1}$ $\propto h(n{+}1)\,\exp\!\bigl(-f\,Q(n{+}1)\bigr)\quad(x_{i}=1).$ (72)
@@
$p(x_{i}=1\mid\mathbf{x}_{\setminus i})\;=\;\frac{1}{1+\exp\!\bigl(-\log(p_{1}/p_{0})\bigr)}.$ (73)
@@
$\log\!\frac{p_{1}}{p_{0}}\;=\;\bigl[\log h(n{+}1)-\log h(n)\bigr]\;-\;f\,\Delta Q(n).$ (74)
For the entropy-canceling base measure $h(n)=1/\binom{N}{n}$,
$\log\!\frac{p_{1}}{p_{0}}\;=\;\log\!\left(\frac{n+1}{N-n}\right)\;-\;f\,\Delta Q(n).$ (75)
@@
The resulting spike sequences were then fitted with the state-space kinetic Ising model. Because the data were stationary, we fixed the state noise covariance to zero, $\mathbf{Q}^{i}=0$ ($i=1,\ldots,N$), and omitted hyperparameter optimization. To reduce computation time, the samples were reorganized into $T=200$ time bins and $R=5000$ trials, preserving dependencies across consecutive bins within each trial. Under this setting, the fitted state-space model yielded constant parameters across bins. We then generated 500,000500{,}000 spike sequences by resampling from the fitted model, and compared their population spike-count distribution with that of the original Gibbs-sampled data.
Supporting information
S1 Fig.
Spike-rate dynamics and distributions for mice 1-18 Spike-rate dynamics and distributions under the active (red) and passive (blue) conditions. The presentation styles for each mouse follow Fig. 6A. The mice were listed in descending order of behavioral performance measured by d-prime. See Fig. S2 Fig for the remaining mice.
S2 Fig.
Spike-rate dynamics and distributions for mice 19-37 The same as in S1 Fig but for the remaining 19 mice.
S3 Fig.
Comparison of mean spike rates and coefficient of variation in the active and passive conditions. A Mean spiking probability across all bins, trials, and neurons in active and passive conditions. Each line represents the same mouse. Neurons showed significantly lower firing rates in the active condition ($p=1.556\times 10^{-8}$, Wilcoxon signed-rank test). B Coefficient of variations (CVs) of the firing rate distributions, a measure of sparseness, in the active and passive conditions. CV was significantly higher in the active condition ($p=8.35\times 10^{-8}$, Wilcoxon signed-rank test).
S4 Fig.
Estimated neural dynamics under active and passive conditions in shuffled data of mouse 574078. Presentation style follows Fig. 6.
S5 Fig.
Mean effective coupling and shuffle control. A Violin plots of time-averaged effective couplings for mouse $574078$ under the active and passive conditions. Horizontal bars indicate the mean (red) and median (green); points show individual matrix entries. B Population summary of the per-mouse mean coupling in the original data; each line connects the active and passive values from the same mouse. Panel annotations report $p$-values and sample size ($n$) from Wilcoxon signed-rank tests across mice (two-sided). C Shuffle-adjusted means, where for each mouse the value in each condition is computed as (Original - Shuffle); lines connect paired values. The annotation reports a one-sided Wilcoxon signed-rank test assessing whether the median of {(Original - Shuffle) in Active} minus {(Original - Shuffle) in Passive} is greater than zero. Together, the results indicate that the mean effective coupling is larger in the active condition than in the passive condition, and that this increase persists after shuffle correction.
S6 Fig.
Time courses of entropy flow and mean spike rates for each mouse under active and passive conditions. Each subplot represents the dynamics of an individual mouse. Solid lines are entropy flows (red for active, blue for passive) while dashed lines represent the average population spike rate (red for active, blue for passive).
S7 Fig.
[…]
S8 Fig.
Relationship between spike rates and entropy flow for an example mouse. A Mean spike rate vs entropy flow per individual unit under the active and passive conditions (mouse 574078). Dashed lines connect values for the two conditions, highlighting behavioral state-dependent changes. B Geometric mean spike rate (abscissa) vs differences in entropy flow (ordinate) for individual units. The positive Spearman correlation coefficient ($\rho=0.22$) suggests that units with higher spike rates increased entropy flow in the active condition.
S9 Fig.
Relating the dependency of entropy flow change of individual units on firing rates with behavioral performance. A (Left) Scatter plot of behavioral performance (mean d-prime) vs. the Spearman rank correlation between the geometric mean rate and entropy flow change of individual units (see Fig. S8 Fig). Each dot represents a single mouse. The dependency in this scatter plot was assessed again by the Spearman rank correlation coefficient, yielding $\rho=0.3578$. (Right) A permutation test comparing the observed correlation value $\rho$ with those obtained from the surrogate data. A statistically significant positive relationship was observed ($p=0.0304$, two-tailed). The surrogate data was constructed by permuting the values of mean d-prime. B Results for trial-shuffled data.
@@
We used publicly available “Allen Brain Observatory: Visual Behavior Neuropixels” dataset set provided by Allen Institute for Brain Science:
@@
The code used to analyze the data is available in the Github repository:

### v3

@@ Methods
${\sigma}_{t}^{\rm forward}$ $\simeq-\sum_{\mathbf{x}_{t-1}}Q(\mathbf{x}_{t-1})$
$\phantom{==}\cdot\sum_{i}\sum_{x_{i,t}}p(x_{i,t}|\mathbf{x}_{t-1})\log p(x_{i,t}|\mathbf{x}_{t-1})$
$=\sum_{i}\sum_{\mathbf{x}_{t-1}}Q(\mathbf{x}_{t-1})\chi\left(h_{i,t}(\mathbf{x}_{t-1})\right),$ (46)
@@
$\phantom{=}-(1-r(h_{i,t}(\mathbf{x}_{t-1})))\log(1-r(h_{i,t}(\mathbf{x}_{t-1})))$
$=-[r(h_{i,t}(\mathbf{x}_{t-1}))h_{i,t}(\mathbf{x}_{t-1})-\psi(h_{i,t}(\mathbf{x}_{t}))]$ (47)
We approximate Eq. 46 by a Gaussian distribution based on the central limit theorem for a collection of independent binary signals. Specifically, by using $\mathcal{D}_{z}=\frac{\mathrm{d}z}{\sqrt{2\pi}}\exp\left(-\frac{1}{2}z^{2}\right)$, the forward conditional entropy is approximated as
${\sigma}_{t}^{\rm forward}$ $\approx\sum_{i}\int\mathcal{D}_{z}\,\chi\left(g_{i,t,t-1}+z\sqrt{\Delta_{i,t,t-1}}\right),$ (48)
where $g_{i,t,t-1}$ and $\Delta_{i,t,t-1}$ are the mean and variance of $h_{i,t}(\mathbf{x}_{t-1})$ given by
$g_{i,t,t-1}$ $=\theta_{i,t}+\sum_{j}\theta_{ij,t}m_{j,t-1},$ (49)
$\Delta_{i,t,t-1}$ $=\sum_{j}\theta_{ij,t}^{2}m_{j,t-1}(1-m_{j,t-1}).$ (50)
Here, $m_{i,t}$ is the mean-field approximation of $x_{i,t}$ obtained by the Gaussian approximation method assuming independent activity of neurons at t1t-1:
$m_{i,t}$ $=\sum_{\mathbf{x}_{t},\mathbf{x}_{t-1}}x_{i,t}\,p(\mathbf{x}_{t}|\mathbf{x}_{t-1})p(\mathbf{x}_{t-1})$
$=\sum_{\mathbf{x}_{t-1}}Q(\mathbf{x}_{t-1})r(h_{i,t}(\mathbf{x}_{t-1})).$ (51)
@@
$m_{i,t}$ $\approx\int\mathcal{D}_{z}\,r\left(g_{i,t,t-1}+z\sqrt{\Delta_{i,t,t-1}}\right),$ (52)
for $t=1,\ldots,T$, using Eqs. 49 and 50, which are functions of $m_{i,t-1}$. Here $m_{i,1}$ was computed using nominal values of $m_{i,0}$ ($i=1,\ldots,N$). In the simulation and empirical analyses, we used spiking probability averaged over all time steps and trials for each neuron as $m_{i,0}$.
@@
$=-\sum_{\mathbf{x}_{t},\mathbf{x}_{t-1}}p(\mathbf{x}_{t}|\mathbf{x}_{t-1})\sum_{\mathbf{x}_{t-2}}p(\mathbf{x}_{t-1}|\mathbf{x}_{t-2})p(\mathbf{x}_{t-2})\log{p(\mathbf{x}_{t-1}|\mathbf{x}_{t})}$
$=-\sum_{\mathbf{x}_{t-2}}\sum_{\mathbf{x}_{t-1}}p(\mathbf{x}_{t-1}|\mathbf{x}_{t-2})p(\mathbf{x}_{t-2})$
$\phantom{==}\cdot\sum_{\mathbf{x}_{t}}p(\mathbf{x}_{t}|\mathbf{x}_{t-1})\sum_{i}\left[x_{i,t-1}h_{i,t}(\mathbf{x}_{t})-\psi(h_{i,t}(\mathbf{x}_{t}))\right].$ (53)
We approximate the following probabilities by independent distributions:
@@
Using them, ${\sigma}^{\rm backward}_{t}$ can be approximated as
${\sigma}^{\rm backward}_{t}\simeq-\sum_{\mathbf{x}_{t-2}}\sum_{\mathbf{x}_{t-1}}p(\mathbf{x}_{t-1}|\mathbf{x}_{t-2})Q(\mathbf{x}_{t-2})$
$\phantom{=}\cdot\sum_{\mathbf{x}_{t}}Q(\mathbf{x}_{t})\sum_{i}\left[x_{i,t-1}h_{i,t}(\mathbf{x}_{t})-\psi(h_{i,t}(\mathbf{x}_{t}))\right]$
$=-\sum_{\mathbf{x}_{t}}Q(\mathbf{x}_{t})\sum_{\mathbf{x}_{t-2}}Q(\mathbf{x}_{t-2})$
$\phantom{=}\cdot\sum_{i}\left[r(h_{i,t-1}(\mathbf{x}_{t-2}))h_{i,t}(\mathbf{x}_{t})-\psi(h_{i,t}(\mathbf{x}_{t}))\right]$
@@
where we used Eq. 43 to obtain the second equality and Eq. 51 to obtain the last result. By defining
$\phi_{i,t}(h_{i,t}(\mathbf{x}_{t}))=-[m_{i,t-1}h_{i,t}(\mathbf{x}_{t})-\psi(h_{i,t}(\mathbf{x}_{t}))],$ (57)
@@
${\sigma}^{\rm backward}_{t}$ $=\sum_{i}\sum_{\mathbf{x}_{t}}Q(\mathbf{x}_{t})\phi_{i,t}(h_{i,t}(\mathbf{x}_{t}))$
$\approx\sum_{i}\int\mathcal{D}_{z}\,\phi_{i,t}\left(g_{i,t,t}+z\sqrt{\Delta_{i,t,t}}\right),$ (58)
@@
$g_{i,t,t}$ $\equiv\theta_{i,t}+\sum_{j}\theta_{ij,t}m_{j,t},$ (59)
$\Delta_{i,t,t}$ $=\sum_{j}\theta_{ij,t}^{2}m_{j,t}(1-m_{j,t}).$ (60)
An alternative approach to obtain the backward conditional entropy is given in Supplementary Note 2.
@@
$\approx\sum_{i}\int\mathcal{D}_{z}\,\Bigg[-\chi\left(g_{i,t,t-1}+z\sqrt{\Delta_{i,t,t-1}}\right)$
$\phantom{==}+\phi_{i,t}\left(g_{i,t,t}+z\sqrt{\Delta_{i,t,t}}\right)\Bigg],$ (61)
@@
See also Supplementary Note 3 for the analytical expression of the entropy flow under steady-state conditions or for independent neurons.
@@
We constructed time-varying field and coupling parameters, from which we generated the binary data. To ensure smooth temporal variations, each coupling parameter $\theta_{ij,t}$ was sampled from a Gaussian process of size $T$ with mean $\mu$ and covariance matrix defined by the squared exponential kernel
$k(t,s)=k_{0}\exp\left({-\frac{(t-s)^{2}}{2\tau^{2}}}\right).$ (62)
For the analysis of estimation error and computational time using different system sizes (Fig. 3), we used the scaling mean $\mu=5/N$ and variance $k_{0}=10/N$, following the convention of the Sherrington-Kirkpatrick model. The characteristic length-scale was specified by $\tau=30/\sqrt{N}$. Similarly, the external field parameters $\theta_{i,t}$ were independently sampled from the Gaussian process, using $\mu=-3$, $\tau=50$, and $k_{0}=1$.
To obtain trajectories for the different system sizes, a single set of random values was generated for the maximum system size, and subsets of these values were used to examine the system size $N$. Specifically, for the coupling parameters, a global three-dimensional array was created with dimensions corresponding to the maximum number of neurons, time steps, and coupling connections. Similarly, for the field parameters, a two-dimensional array was generated, with dimensions corresponding to the maximum number of time steps and neurons. For a given neuron count $N$, the relevant subset of values was extracted from these precomputed arrays, ensuring that each $N$ used a subset of the values assigned to larger $N$. This hierarchical structure ensured that the seed for $N=80$ encompassed all values used for smaller $N$, maintaining consistency across different system sizes. We evaluated the model’s performance using this data set and repeated the procedure 10 times.
@@
To perform the analysis on fitting the kinetic Ising model to a mismatched model, we generated binary spike sequences using a nonlinearity that goes beyond linear synaptic summation and a logistic activation function, which therefore induces the higher-order interactions (HOIs) in the population activity. For this goal, we employed the recently proposed alternating-shrinking HOI model [58].
@@
$p(\mathbf{x})=\frac{h\left(\sum_{i=1}^{N}x_{i}\right)}{Z}\exp\left[-f\sum_{j=1}^{N}\left(-1\right)^{j+1}C_{j}\left(\frac{\sum_{i}x_{i}}{N}\right)^{j}\right],$ (63)
where $f$ is a sparsity parameter and $Z$ is the partition function. Let $n=\sum_{i=1}^{N}x_{i}$. $h\left(n\right)$ is an entropy-canceling base measure function defined using the binomial coefficient:
$h\left(n\right)=1\left/\binom{N}{n}\right$.. (64)
The parameters $C_{1},C_{2},\ldots,C_{N}$ are the shrinking parameters, where $C_{j}=\left(\tau\right)^{j}$ with  $0<\tau<1$ results in the shifted-geometric population spike-count distribution.
The population spike-count distribution is the probability distribution of $n$ active neurons in the binary patterns, which is given as
$P\left(n\right)=$ $\binom{N}{n}p\left(x_{1}=1,\ldots,x_{n}=1,x_{n+1}=0,\ldots,x_{N}=0\right)$
@@
We performed Gibbs sampling from this distribution, which dictates the dynamics of a recurrent neural network with threshold-supralinear activation nonlinearity. For neuron $i$, let $\tilde{n}=\sum_{j\neq i}x_{j}$ be the spike count of the other units, and define
[…]
The unnormalized joint activities of neurons are
[…]
$p_{1}$ $\propto h(\tilde{n}{+}1)\,\exp\!\bigl(-f\,Q(\tilde{n}{+}1)\bigr)\quad$ $(x_{i}=1).$ (69)
@@
$p(x_{i}=1{|}\mathbf{x}_{\setminus i})\;=\;\frac{1}{1+\exp\!\bigl(-\log(p_{1}/p_{0})\bigr)}.$ (70)
@@
[…]
@@
The resulting spike sequences were then fitted with the state-space kinetic Ising model. Because the data were stationary, we fixed the state noise covariance to zero, $\mathbf{Q}^{i}=0$ ($i=1,\ldots,N$), and omitted hyperparameter optimization. To reduce computation time, the samples were reorganized into $T=200$ time bins and $L=5000$ trials, preserving dependencies across consecutive bins within each trial. Under this setting, the fitted state-space model yielded constant parameters across bins. We then generated 500,000500{,}000 spike sequences by resampling from the fitted model, and compared their population spike-count distribution with that of the original Gibbs-sampled data.
@@
We used the publicly available Allen Brain Observatory: Visual Behavior Neuropixels dataset provided by the Allen Institute for Brain Science:
@@
Large precomputed datasets required to reproduce the figures are archived on Zenodo:
doi:10.5281/zenodo.15220108.
@@
The analysis code used in this study is archived on Zenodo and linked to the GitHub repository:
doi:10.5281/zenodo.17504162.
For convenient browsing, see the GitHub mirror: